# Efficient Bayesian Additive Regression Models For Microbiome Studies

### v1

@@ 2.1 Defining Marginally Latent Matrix-T Processes (MLTPs)
where $g$ is any distribution depending on parameters $\mathbf{\Pi}$ as well as hyperparameters $\lambda$ and $\phi$ is a known transform. LTPs can alternatively be written as a joint model $p(\mathbf{Y},\mathbf{H})$.
A stochastic process $\mathbf{Y}$ is Marginally LTP (MLTP) if it can be described by a joint distribution $p(\mathbf{Y},\mathbf{H},\mathbf{\Phi})$ with a marginal $p(\mathbf{Y},\mathbf{H})$ that is a LTP. Silverman et al. (2022) showed that a wide variety of linear, dynamic linear, and non-linear regression models are MLTP. In Section 3, we show that our proposed class of generalized additive Gaussian process regression models are MLTP as well.
2.2 Bayesian Multinomial Logistic Normal MLTPs (MLN-MLTPs)
Bayesian Multinomial Logistic Normal MLTPs (MLN-MLTPs) are a subtype of MLTPs that are particularly useful for the analysis of microbiome data. In MLN-MLTPs, the distribution $g$ is a product multinomial: $p(\mathbf{Y}_{\cdot 1},\dots,\mathbf{Y}_{\cdot N})\sim\prod_{n=1}^{N}\text{Multinomial}(\boldsymbol{\Pi}_{\cdot n})$ and the transform $\phi$ is an invertible log-ratio transform from the $D$-dimensional simplex to $D-1$ dimentional real-space: $\mathbf{H}_{\cdot n}=\phi(\boldsymbol{\Pi}_{\cdot n}\in\mathbb{S}^{D})\in\mathbb{R}^{D-1}$. Canonically, we used the following Additive Log-Ratio (ALR) transform which takes the $D$-th taxa as a reference:
$\mathbf{H}_{.n}=\phi(\boldsymbol{\Pi}_{.n})=\left\{\log\left(\frac{\pi_{1n}}{\pi_{Dn}}\right),\ldots,\log\left(\frac{\pi_{(D-1)n}}{\pi_{Dn}}\right)\right\}^{T}.$ (1)
We choose this transform for computational efficiency as discussed in Silverman et al. (2022). There is no loss in generality as posterior samples taken with respect to the $\text{ALR}_{D}$ coordinate system can be transformed into any other log-ratio coordinate system (Pawlowsky-Glahn, Egozcue and Tolosana-Delgado, 2015, Appendix A.3). For context, the $\text{ALR}_{D}$ transform is the inverse of the softmax transform.
2.3 Collapsed-Uncollapsed (CU) Sampler
The definition of MLTPs is key to efficient inference. If a model $p(\mathbf{Y},\mathbf{H},\mathbf{\Phi})$ has a closed-form marginal $p(\mathbf{Y},\mathbf{H})$ that is an LTP, then its closed form conditional $p(\mathbf{\Phi}\mid\mathbf{Y},\mathbf{H})$ likely exists. We call the marginal $p(\mathbf{Y},\mathbf{H})$ the collapsed form and $p(\mathbf{\Phi}\mid\mathbf{Y},\mathbf{H})$ the uncollapsed form. The posterior of an MLTP factors as
$p(\mathbf{H},\Phi\mid\mathbf{Y})=p(\mathbf{\Phi}\mid\mathbf{H},\mathbf{Y})p(\mathbf{H}\mid\mathbf{Y})$
with the uncollapsed form as the first term and the posterior of the collapsed form as the second. As the collapsed form is rarely conjugate, techniques such as MCMC can be used to obtain samples from it’s posterior. Then, conditioned on those samples, the uncollapsed form can be used to obtain samples from the joint posterior. Especially when $\boldsymbol{\Phi}$ is high-dimensional, this Collapse-Uncollapse sampler can be much more efficient than common alternatives (Silverman et al., 2022). Still, the most substantial enhancements occur when approximations to the collapsed form are considered.
We have developed a Laplace approximation for the collapsed form of MLTPs:
$p(vec(\mathbf{H})\mid\mathbf{Y})\approx N(vec(\hat{\mathbf{H}}),\nabla^{-2}[vec(\hat{\mathbf{H}})])$
where $\hat{\mathbf{H}}$ denotes the Maximum A Posteriori (MAP) estimate of the collapsed form and $\nabla^{-2}[vec(\hat{\mathbf{H}})]$ denotes the inverse Hessian of the collapsed form evaluated at the MAP estimate. To facilitate this approximation, we derived analytical results for the gradient and Hessian of the collapsed form (Silverman et al., 2022). Focusing on applications to MLN-MLTPs, we proved error bounds on the Laplace approximation and provided simulation and real analyses showing that the approximation was extremely accurate in the context of microbiome data analysis. Beyond the accuracy of posterior calculations, we showed that this CU sampler with the Laplace approximation (simply referred to as the CU sampler in the following text) was often 4-5 orders of magnitude faster than MCMC and 1-2 orders of magnitude faster than black-box variational inference while also being more accurate than the later. The CU sampler for MLN-MLTPs, along with uncollapse samplers for linear and non-linear regression models, is publicly available on CRAN as part of the fido software package.
3 Methods
3.1 Multinomial Logistic Normal Additive Gaussian Process Models (MultiAddGP)
To facilitate additive linear and non-linear modeling using Bayesian MLNs, this article introduces the class of Multinomial Logistic Normal Additive Gaussian Process Models (MultiAddGPs). Let $\mathbf{Y}_{\cdot n}$ denote a $D$-vector of observed data, $\mathbf{X}_{\cdot n}$ a $Q_{0}$-vector of covariates to model linearly, and each $\mathbf{Z}^{(k\in\{1,\dots,K\})}_{\cdot n}$ a $Q_{k}$-vector of covariates to be modeled with distinct non-linear functions. MultiAddGP models have the following form:
$\boldsymbol{\Pi}_{\cdot n}$ $=\phi^{-1}(\mathbf{H}_{\cdot n})$ (3)
$\mathbf{H}_{\cdot n}$ $\sim N(\mathbf{F}_{\cdot n},\boldsymbol{\Sigma})$ (4)
with priors $\mathbf{B}\sim N(\boldsymbol{\Theta}^{(0)},\boldsymbol{\Sigma},\boldsymbol{\Gamma}^{(0)})$, $\mathbf{f}^{(k)}\sim\text{GP}(\boldsymbol{\Theta}^{(k)},\boldsymbol{\Sigma},\boldsymbol{\Gamma}^{(k)})$, and $\boldsymbol{\Sigma}\sim\text{InvWishart}(\boldsymbol{\Xi},\zeta)$. As in prior sections, $\phi$ denotes any log-ratio transform from $\mathbb{S}^{D}$ to $\mathbb{R}^{D-1}$. $\boldsymbol{\Sigma}$ is a $(D-1)\times(D-1)$ covariance matrix. For the matrix-normal prior on the linear term, $\boldsymbol{\Theta}^{(0)}$ is the mean matrix and $\boldsymbol{\Gamma}^{(0)}$ is a $Q_{0}\times Q_{0}$ covariance matrix representing covariance in the parameters of the $Q_{0}$ covariates. The terms $\boldsymbol{\Theta}^{(k)}$ and $\boldsymbol{\Gamma}^{(k)}$ in the $K$ matrix-normal process priors echo their linear counterparts but are functions (e.g., mean and kernel functions) rather than fixed dimensional matrices. Additionally, we permit hyperparameters $\Omega$ in any part of the model including but not limited to hyperparameters in the kernel functions (e.g., bandwidth parameters) or hyperparameters in the mean functions. As we will show through simulated and real data analyses in Sections 4 and 5, this is a very flexible form of model which can be used in a wide range of additive linear and non-linear modeling tasks.
3.2 Posterior Estimation in MultiAddGPs
We use MLTP theory to sample from the posterior of MultiAddGPs: $p(\mathbf{H},\mathbf{B},\mathbf{f}^{(1)},\dots,\mathbf{f}^{(K)},\boldsymbol{\Sigma}\mid\mathbf{Y})$. To connect to that theory we first prove that MultiAddGPs are MLN-MLTP’s with $\boldsymbol{\Phi}=\{\mathbf{B},\mathbf{f}^{(1)},\dots,\mathbf{f}^{(K)},\boldsymbol{\Sigma}\}$. We derive the collapsed form and then derive an efficient algorithm for sampling from the uncollapsed form. First however, we must clarify the definition of $\mathbf{F},\mathbf{f}_{1},\dots,\mathbf{f}_{K}$.
@@ 3.2 Posterior Estimation in MultiAddGPs
Up to this point, we have not distinguished between the set of points $n\in\{1,\dots,N\}$ at which we have observed data $\mathbf{Y}$ and the potentially different set $n^{*}\in\{1,\dots,N^{*}\}$ at which we want to evaluate the functions $\mathbf{F},\mathbf{f}_{1},\dots,\mathbf{f}_{K}$. In what follows, we use the symbols $\mathbf{F},\mathbf{f}_{1},\dots,\text{ and }\mathbf{f}_{K}$ to denote the evaluation of corresponding infinite-dimensional functions at the set of evaluation points $\{1,\dots,N^{*}\}$, i.e., they are each $D\times N^{*}$-dimensional random matrices. In contrast, all other random matrices (e.g., $\mathbf{H}$) represent their corresponding infinite-dimensional analogues evaluated at the set of observed points ($\{1,\dots,N\}$).
In Supplementary Section 1 we prove that MultiAddGPs are MLN-MLTPs with the following parameters defining the collapsed form:
$\mathbf{M}$ $=\boldsymbol{\Theta}^{(0)}\mathbf{X}+\sum_{k=1}^{K}\boldsymbol{\Theta}^{(k)}(\mathbf{Z}^{(k)})$
$\mathbf{V}$ $=\boldsymbol{\Xi}$
$\mathbf{A}$ $=\mathbf{X}^{T}\boldsymbol{\Gamma}^{(0)}\mathbf{X}+\sum_{k=1}^{K}\boldsymbol{\Gamma}^{(k)}(\mathbf{Z}^{(k)})+\mathbf{I}$
$\mathbf{\nu}$ $=\mathbf{\zeta}$
[…]
@@
Sampling from the uncollapsed form starts by obtaining samples from $p(\mathbf{F},\boldsymbol{\Sigma}\mid\mathbf{H})$. Conditioning on $\mathbf{H}$ and marginalizing over $\mathbf{f}^{(1)},\dots,\mathbf{f}^{(K)}$ in the MultiAddGP model results in a Bayesian matrix-normal process model with likelihood $\mathbf{H}_{\cdot n}\sim N(\mathbf{F}_{\cdot n},\boldsymbol{\Sigma})$ and priors:
@@
$\mathbf{F}$ $\sim\text{GP}\left(\boldsymbol{\Theta}^{(0)}\mathbf{X}+\sum_{k=1}^{K}\boldsymbol{\Theta}^{(k)}(\mathbf{Z}^{(k)}),\;\boldsymbol{\Sigma},\;\mathbf{X}^{T}\boldsymbol{\Gamma}^{(0)}\mathbf{X}+\sum_{k=1}^{K}\boldsymbol{\Gamma}^{(k)}(\mathbf{Z}^{(k)})\right)$
$\boldsymbol{\Sigma}$ $\sim\text{InvWishart}(\boldsymbol{\Xi},\zeta).$
This is the same model discussed in Silverman et al. (2022): samples from $p(\mathbf{F},\boldsymbol{\Sigma}\mid\mathbf{H})$ can be obtained via methods described in Appendix C of that article.
Finally, conditioned on samples of $\mathbf{F}$ and $\boldsymbol{\Sigma}$, we obtain samples of each $\mathbf{f}^{(k)}$. Inspired by the backfitting algorithm used for estimation in generalized additive models (Hastie and Tibshirani, 1990), we developed a backsampling algorithm for this task. The backsampling proceeds by iteratively sampling $p(\mathbf{B}\mid\mathbf{F})$, $p(\mathbf{f}_{1}\mid\mathbf{F},\mathbf{B})$, $p(\mathbf{f}_{2}\mid\mathbf{F},\mathbf{B},\mathbf{f}_{1})$, $\dots$, and then $p(\mathbf{f}^{(K)}\mid\mathbf{F},\mathbf{B},\mathbf{f}_{1},\dots,\mathbf{f}^{(K-1)})$. For brevity, we leave a description of this algorithm to Supplementary Sections 2 and 3.

### v2

@@ 2.2.1 MultiAddGPs as MLTPs and the Collapse Step
where $g$ is any distribution depending on parameters $\mathbf{\Pi}$ and $\phi$ is a known transform.
A stochastic process $\mathbf{Y}$ is Marginally LTP (MLTP) if there exists a joint distribution $p(\mathbf{Y},\mathbf{H},\mathbf{\Phi})$ such that the marginal $p(\mathbf{Y},\mathbf{H})$ is a LTP. In Supplementary Section 2, we prove the following theorem, which establishes that MultiAddGPs are MLTPs.
Theorem 2.3.
MultiAddGPs are MLTPs with parameters $\boldsymbol{\Phi}=\{\mathbf{F},\mathbf{B},\mathbf{f}^{(1)},\dots,\mathbf{f}^{(K)},\boldsymbol{\Sigma}\}$. The collapsed form $p(\mathbf{Y},\mathbf{H})$ is an LTP:
$\mathbf{Y}_{\cdot n}$ $\sim\text{Multinomial}(\mathbf{\Pi}_{\cdot n})$
$\mathbf{\Pi}_{\cdot n}$ $=\phi^{-1}(\mathbf{H}_{\cdot n})$
$\mathbf{H}$ $\sim TP(\nu,\mathbf{M},\mathbf{V},\mathbf{A})$
where $\phi$ is an invertible log-ratio transformation, $\mathbf{M}=\boldsymbol{\Theta}^{(0)}\mathbf{X}+\sum_{k=1}^{K}\boldsymbol{\Theta}^{(k)}(\mathbf{Z}^{(k)})$, $\mathbf{V}=\boldsymbol{\Xi}$, and $\mathbf{A}=\mathbf{X}^{T}\boldsymbol{\Gamma}^{(0)}\mathbf{X}+\sum_{k=1}^{K}\boldsymbol{\Gamma}^{(k)}(\mathbf{Z}^{(k)})+\mathbf{I}_{N},\nu=\zeta$
In the collapsed step of the CU sampler, we approximate $p(\mathbf{H}\mid\mathbf{Y})$ using a Laplace approximation: $q(\mathbf{H}|\mathbf{Y})=\mathcal{N}\left(\text{vec}(\hat{\mathbf{H}}),\nabla^{-2}(\text{vec}(\hat{\mathbf{H}}))\right)$, where $\hat{\mathbf{H}}$ is the maximum a posteriori (MAP) estimate of $p(\mathbf{H}\mid\mathbf{Y})$, and $\nabla^{-2}(\text{vec}(\hat{\mathbf{H}}))$ is the inverse Hessian of $\log p(\mathbf{H}\mid\mathbf{Y})$ evaluated at $\hat{\mathbf{H}}$.
We obtain efficient MAP estimates $\hat{\mathbf{H}}$ by applying the L-BFGS optimizer to the closed-form gradients $\nabla\bigl{(}\mathrm{vec}(\hat{\mathbf{H}})\bigr{)}$ derived in Supplementary Section 3. Building on these results, we then compute the Hessian $\nabla^{-2}\bigl{(}\mathrm{vec}(\hat{\mathbf{H}})\bigr{)}$ in closed form (also see Supplementary Section 3). This approximation is computationally efficient and extremely accurate. The error in this approximation is bounded in probability by $O_{p}([D-1]\sum_{n=1}^{N}\lambda_{n}^{-1})$ where $\lambda_{n}=\sum_{d=1}^{D}\mathbf{Y}_{dn}$ (See Proposition 11 of Silverman et al. (2022)). This bound explains the high accuracy of the approximation in sequence count applications, such as microbiome data analysis, where the sequencing depth $\lambda_{n}$ is typically large.
2.2.2 Back-sampling and the Uncollapse Step for MultiAddGPs
Each posterior sample from $p(\mathbf{H},\mid\mathbf{Y})$ is then conditioned upon and used to sample from $p(\mathbf{\Phi}\mid\mathbf{H},\mathbf{Y})$ where $\boldsymbol{\Phi}=\{\mathbf{F},\mathbf{B},\mathbf{f}^{(1)},\dots,\mathbf{f}^{(K)},\boldsymbol{\Sigma}\}$. Since $\mathbf{\Phi}$ is conditionally independent of $\mathbf{Y}$ given $\mathbf{H}$ ($\mathbf{\Phi}\perp\mathbf{Y}\mid\mathbf{H}$), the uncollapse step simplifies to deriving an efficient sampler for $p(\mathbf{\Phi}\mid\mathbf{H})$. Further, using the fact that $\{\mathbf{B},\mathbf{f}^{(1)},\dots,\mathbf{f}^{(K)}\}\perp\mathbf{H}\mid\{\mathbf{F},\boldsymbol{\Sigma}\}$ we can break the uncollapse step into two simpler steps, with the second conditioned on the results of the first: (1) sample $p(\boldsymbol{\Sigma},\mathbf{F}\mid\mathbf{H})$ and (2) sample $p(\mathbf{B},\mathbf{f}^{(1)},\dots,\mathbf{f}^{(K)}\mid\mathbf{F},\boldsymbol{\Sigma})$.
@@ 2.2.2 Back-sampling and the Uncollapse Step for MultiAddGPs
Before proceeding, we must clarify the distinction between the set of observed data points $n\in\{1,\dots,N\}$, where the response $\mathbf{Y}$ is observed, and the evaluation set $n^{*}\in\{1,\dots,N^{*}\}$, where the functions $\mathbf{F},\mathbf{B}_{0},\mathbf{f}^{(1)},\dots,\mathbf{f}^{(K)}$ are to be inferred. In what follows, we use the symbols $\mathbf{F},\mathbf{B}_{0},\mathbf{f}^{(1)},\dots,\mathbf{f}^{(K)}$ to denote the evaluation of corresponding infinite-dimensional functions at the set of evaluation points $\{1,\dots,N^{*}\}$, i.e., they are each $D\times N^{*}$-dimensional random matrices. In contrast, all other random matrices (e.g., $\mathbf{H}$) represent their corresponding infinite-dimensional analogues evaluated at the set of observed points $\{1,\dots,N\}$.
2.2.2.1 Sampling $p(\boldsymbol{\Sigma},\mathbf{F}\mid\mathbf{H})$
@@
Conditioning on $\mathbf{H}$ and marginalizing over $\mathbf{f}^{(k)},k\in\{1,\dots,K\}$ in the MultiAddGPs model results in a Bayesian matrix-normal process model with likelihood $\mathbf{H}_{\cdot n}\sim\mathcal{N}(\mathbf{F}_{\cdot n},\boldsymbol{\Sigma})$ and priors:
@@
$\mathbf{F}$ $\sim\mathcal{GP}\left(\boldsymbol{\Theta}^{(0)}\mathbf{X}+\sum_{k=1}^{K}\boldsymbol{\Theta}^{(k)}(\mathbf{Z}^{(k)}),\;\boldsymbol{\Sigma},\;\mathbf{X}^{T}\boldsymbol{\Gamma}^{(0)}\mathbf{X}+\sum_{k=1}^{K}\boldsymbol{\Gamma}^{(k)}(\mathbf{Z}^{(k)})\right)$
$\boldsymbol{\Sigma}$ $\sim\mathcal{IW}(\boldsymbol{\Xi},\zeta).$
This model is conjugate and posterior samples of $\mathbf{F}$ and $\mathbf{\Sigma}$ can be efficiently obtained as described in Supplementary Section 2.2.
2.2.2.2 Sampling $p(\mathbf{B},\mathbf{f}^{(1)},\dots,\mathbf{f}^{(K)}\mid\mathbf{F},\boldsymbol{\Sigma})$ via Back-sampling
Inspired by the backfitting algorithm used in generalized additive models (Wood 2017), we introduce a novel Back-sampling (BS) algorithm which iteratively smooths (i.e., decomposes) $\mathbf{F}$ into additive components.
For each posterior sample, the BS algorithm proceeds by sampling $p(\mathbf{B}\mid\mathbf{F},\boldsymbol{\Sigma})$ and then sequentially sampling $p(\mathbf{f}^{(1)}\mid\mathbf{B},\mathbf{F},\boldsymbol{\Sigma}),\dots,p(\mathbf{f}^{(K-1)}\mid\mathbf{f}^{(1)},...,\mathbf{f}^{(K-2)},\mathbf{B},\mathbf{F},\boldsymbol{\Sigma})$. Note the value of the final $\mathbf{f}^{(K)}$ is uniquely determined by the values of $\mathbf{F},\mathbf{B},\mathbf{f}^{(1)},\dots,\mathbf{f}^{(K-1)}$ and so a sample of $\mathbf{f}^{(K)}$ from the joint distribution can be obtained by calculating $\mathbf{f}^{(K)}=\mathbf{F}-\mathbf{B}\mathbf{X}-\mathbf{f}^{(1)}-\cdots-\mathbf{f}^{(K-1)}$. The pseudocode for the BS algorithm is provided in Algorithm 2.
First we sample from the posterior $p(\mathbf{B}\mid\mathbf{F},\boldsymbol{\Sigma})$. To do this we marginalize over $\mathbf{f}^{(1)},\dots,\mathbf{f}^{(K)}$ and reparameterize as $\mathbf{B}^{\epsilon}:=\mathbf{F}-\sum_{k=1}^{K}\mathbf{\Theta}^{(k)}(\mathbf{Z}^{(k)})$.
This leads to the following conjugate Bayesian model for $\mathbf{B}$:
$\mathbf{B}^{\epsilon}$ $\sim\mathcal{MN}(\mathbf{BX},\boldsymbol{\Sigma},\boldsymbol{\Gamma}^{*})$
$\mathbf{B}$ $\sim\mathcal{MN}(\boldsymbol{\Theta}^{(0)},\boldsymbol{\Sigma},\boldsymbol{\Gamma}^{(0)})$
where $\boldsymbol{\Gamma}^{*}=\sum_{k=1}^{K}\boldsymbol{\Gamma}^{(k)}(\mathbf{Z}^{(k)})$, and $\boldsymbol{\Theta}^{(0)}$, $\boldsymbol{\Gamma}^{(0)}$ denote the prior mean and covariance of $\mathbf{B}$, respectively. Using Supplementary Section 4, we have the following closed form for the posterior:
$\mathbf{B}\mid\mathbf{B}^{\epsilon},\boldsymbol{\Sigma}\sim\mathcal{MN}\Bigg{(}\left[\mathbf{B}^{\epsilon}\boldsymbol{\Gamma}^{-*}\mathbf{X}^{T}+\boldsymbol{\Theta}^{(0)}\boldsymbol{\Gamma}^{-(0)}\right]\left[\mathbf{X}\boldsymbol{\Gamma}^{-*}\mathbf{X}^{T}+\boldsymbol{\Gamma}^{-(0)}\right]^{-1},\boldsymbol{\Sigma},\;\left[\mathbf{X}\boldsymbol{\Gamma}^{-*}\mathbf{X}^{T}+\boldsymbol{\Gamma}^{-(0)}\right]^{-1}\Bigg{)}$ where $\boldsymbol{\Gamma}^{-*}=(\boldsymbol{\Gamma}^{*})^{-1}$ and $\boldsymbol{\Gamma}^{-(0)}=(\boldsymbol{\Gamma}^{(0)})^{-1}$.
We use a similar trick to sample each nonlinear component $p(\mathbf{f}^{(1)}\mid\mathbf{B},\mathbf{F},\boldsymbol{\Sigma}),\dots,p(\mathbf{f}^{(K-1)}\mid\mathbf{f}^{(1)},...,\mathbf{f}^{(K-2)},\mathbf{B},\mathbf{F},\boldsymbol{\Sigma})$. For each $k\in\{1,\dots,K-1\}$, we reparameterize the model as: $\mathbf{f}^{\epsilon_{k}}=\mathbf{F-BX}-\sum_{i=1}^{k-1}\mathbf{f}^{(i)}-\sum_{j=k+1}^{K}\boldsymbol{\Theta}^{(j)}(\mathbf{Z}^{(j)})$. Similar to $\mathbf{B}^{\epsilon}$, $\mathbf{f}^{\epsilon_{k}}$ leads to the following conjugate Bayesian model for each $\mathbf{f}^{(k)}$:
$\mathbf{f}^{\epsilon_{k}}$ $\sim\mathcal{MN}(\mathbf{f}^{(k)},\boldsymbol{\Sigma},\boldsymbol{\Gamma}^{*})$
$\mathbf{f}^{(k)}$ $\sim\mathcal{GP}(\boldsymbol{\Theta}^{(k)},\boldsymbol{\Sigma},\boldsymbol{\Gamma}^{(k)})$
where $\boldsymbol{\Gamma}^{*}=\sum_{j=k+1}^{K}\boldsymbol{\Gamma}^{(j)}$. Again, this can be sampled in closed form using the results of Supplementary Section 4:
$\mathbf{f}^{(k)}\mid\boldsymbol{\Sigma},\mathbf{f}^{\epsilon_{k}}\sim\mathcal{MN}\Bigg{(}\left[\mathbf{f}^{\epsilon_{k}}\boldsymbol{\Sigma}^{-*}+\boldsymbol{\Theta}^{(k)}\boldsymbol{\Gamma}^{-(k)}\right]\left[\boldsymbol{\Gamma}^{-*}+\boldsymbol{\Gamma}^{-(k)}\right]^{-1},\boldsymbol{\Sigma},\left[\boldsymbol{\Gamma}^{-*}+\boldsymbol{\Gamma}^{-(k)}\right]^{-1}\Bigg{)}$ where $\boldsymbol{\Gamma}^{-*}$ and $\boldsymbol{\Gamma}^{-(k)}$ are short-hand for $(\boldsymbol{\Gamma}^{*})^{-1}$ and $(\boldsymbol{\Gamma}^{(k)})^{-1}$ respectively.
Finally, the last component $\mathbf{f}^{(K)}$ is uniquely determined as: $\mathbf{f}^{(K)}=\mathbf{F}-\mathbf{BX}-\sum_{k=1}^{K-1}\mathbf{f}^{(k)}$
Algorithm 2 Back-sampling (BS)
Input: $\{\mathbf{Y,X,Z}\}$ are data observations; $\{\mathbf{F},\boldsymbol{\Sigma}\}$ are samples from the uncollapsed step (1); $\Delta=\{\mathbf{\Theta}^{(0)},\dots,\mathbf{\Theta}^{(K)},\mathbf{\Gamma}^{(0)},\dots,\mathbf{\Gamma}^{(K)}\}$ is a set of prior inputs.
Output: $S$ samples of the form $\{\mathbf{B},\mathbf{f}^{(1)},\dots,\mathbf{f}^{(K)}\}$
1:for $s=1$ to $S$ do
2:     $\mathbf{B}^{\epsilon}=\mathbf{F}-\sum_{k=1}^{K}\boldsymbol{\Theta}^{(k)}(\mathbf{Z}^{(k)})$
3:     $\boldsymbol{\Gamma}^{*}=\sum_{k=1}^{K}\boldsymbol{\Gamma}^{(k)}$
4:     $\mathbf{B}|\mathbf{B}^{\epsilon},\boldsymbol{\Sigma}\sim\mathcal{MN}\Bigg{(}\left[\mathbf{B}^{\epsilon}\boldsymbol{\Gamma}^{-*}\mathbf{X}^{T}+\boldsymbol{\Theta}^{(0)}\boldsymbol{\Gamma}^{-(0)}\right]\left[\mathbf{X}\boldsymbol{\Gamma}^{-*}\mathbf{X}^{T}+\boldsymbol{\Gamma}^{-(0)}\right]^{-1},\;\boldsymbol{\Sigma},\;\left[\mathbf{X}\boldsymbol{\Gamma}^{-*}\mathbf{X}^{T}+\boldsymbol{\Gamma}^{-(0)}\right]^{-1}\Bigg{)}$
5:     for $k=1$ to $K$ do
6:         $\mathbf{f}^{\epsilon_{k}}=\mathbf{F-BX}-\sum_{i=1}^{k-1}\mathbf{f}^{(i)}-\sum_{j=k+1}^{K}\boldsymbol{\Theta}^{(j)}(\mathbf{Z}^{(j)})$
7:         $\boldsymbol{\Gamma}^{*}=\sum_{j=k+1}^{K}\boldsymbol{\Gamma}^{(j)}$
8:         $\mathbf{f}^{(k)}\mid\boldsymbol{\Sigma},\mathbf{f}^{\epsilon_{k}}\sim\mathcal{MN}\left(\left[\mathbf{f}^{\epsilon_{k}}\boldsymbol{\Sigma}^{-*}+\boldsymbol{\Theta}^{(k)}\boldsymbol{\Gamma}^{-(k)}\right]\left[\boldsymbol{\Gamma}^{-*}+\boldsymbol{\Gamma}^{-(k)}\right]^{-1},\;\boldsymbol{\Sigma},\;\left[\boldsymbol{\Gamma}^{-*}+\boldsymbol{\Gamma}^{-(k)}\right]^{-1}\right)$
9:     end for
10:     $\mathbf{f}^{(K)}=\mathbf{F}-\mathbf{BX}-\sum_{k=1}^{K-1}\mathbf{f}^{(k)}$
11:end for